OpenStudy (anonymous):

Solve the equation: 4x2 = 9 {-2, -3} {2/3, -2/3} {3/2, -3/2} no solution

OpenStudy (alexwee123):

{3/2, -3/2}

OpenStudy (unklerhaukus):

\[4x^2 = 9\] \[x^2=\frac{9}{4}\] \[±x=3/2\]